Anders Company provided the following information:

Turnover

1.4

Average operating assets

$120,000

Operating income

$6,720

What is margin?

c.Cannot be determined from the above information.

A company had WACC (weighted average cost of capital) equal to 8.96% If the company pays off mortgage bonds with an interest rate of 4% and issues an equal amount of new stock considered to be relatively risky by the market, which of the following is true?

a.residual income will increase.

b.WACC will decrease.

c.ROI will decrease.

d.WACC will increase.
Answer #1

1. Sales = turnover*average operating assets

= 1.4*120000 = $168000

Margin operating income / sales

= 6720/ 168000

= 4%

2. WACC will increase.